OpenStudy (anonymous):

Write the equation of the line passing through the point (3, 2) and perpendicular to the line .

OpenStudy (anonymous):

y=1/4x + 4

OpenStudy (anonymous):

k*k_1=-1 substitute... find k ... then we'll see.

OpenStudy (anonymous):

slope of the line perpendicular to your equation = -1 / (1/4) = -4 now use the general form below to insert x1 = 3 and y1 = 2 and m = -4 and then simplify y-y1 = m(x - x1)

OpenStudy (dream95):

y-y1=m(x-x1) use the formula m=gradient

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

hmmm>???!!!

OpenStudy (anonymous):

thinking

OpenStudy (dream95):

the gradient is perpendicular to the line so m= -4

OpenStudy (anonymous):

@DuaneC just follow the steps that joywhite told you...

OpenStudy (anonymous):

ok i got it thx